Mill Creek Corporation (MCC) currently has 475,000 shares of stock outstanding that sell for $80 per share. Assuming no market imperfections or tax effects exist, what will the share price be after:

a. MCC has a four-for-three stock split? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price ?

b. MCC has a 20% stock dividend? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price ?

c. MCC has a 52.5% stock dividend? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price ?

d. MCC has a three-for-seven reverse stock split? (Do not round intermediate calculations. Round the final answer to 2 decimal places. Omit $ sign in your response.)

New share price ?

Determine the new number of shares outstanding in parts (a) through (d). (Do not round intermediate calculations. Round the final answers to the nearest whole number.)
Answer #1

1.
=80/4*3=60.00

2.
=80/1.2=66.6666666666667

3.
=80/1.525=52.4590163934426

4.
=80/7*3=34.2857142857143

1.
=475000/3*4=633333.333333333

2.
=475000*1.2=570000.00

3.
=475000*1.525=724375.00

4.
=475000*3/7=203571.428571429